# Automatic Target-Less Camera-LiDAR Calibration From Motion and Deep Point Correspondences

### v1

@@ Abstract
Sensor setups of robotic platforms commonly include both camera and LiDAR as they provide complementary information. However, fusing these two modalities typically requires a highly accurate calibration between them. In this paper, we propose MDPCalib which is a novel method for camera-LiDAR calibration that requires neither human supervision nor any specific target objects. Instead, we utilize sensor motion estimates from visual and LiDAR odometry as well as deep learning-based 2D-pixel-to-3D-point correspondences that are obtained without in-domain retraining. We represent the camera-LiDAR calibration as a graph optimization problem and minimize the costs induced by constraints from sensor motion and point correspondences. In extensive experiments, we demonstrate that our approach yields highly accurate extrinsic calibration parameters and is robust to random initialization. Additionally, our approach generalizes to a wide range of sensor setups, which we demonstrate by employing it on various robotic platforms including a self-driving perception car, a quadruped robot, and a UAV. To make our calibration method publicly accessible, we release the code on our project website at http://calibration.cs.uni-freiburg.de.
I Introduction
Sensor fusion for robotic systems has been extensively investigated [1, 2] as it promises to efficiently combine complementary information from different modalities, e.g., to increase robustness in case of sensor failures [3] and towards weather conditions [4]. However, the effectiveness of fusion approaches heavily depends on the extrinsic calibration between the sensors such as cameras and LiDAR.
Due to the importance of the task, camera-LiDAR calibration has been widely studied by the research community. Previously proposed methods can generally be classified into target-based and target-less approaches. Approaches of the first category often rely on artificial patterns such as checkerboards [5, 6] and require manual labor or another kind of human supervision to associate 2D points from the image space with 3D points in the LiDAR point cloud [7]. While a substantial effort has gone into the detection of the target and the automation of the matching process [8, 9], calibration often still needs special data collection. Some target-less calibration methods aim at overcoming this problem, e.g., by inferring the extrinsic transform from the sensor motion [10] or by matching vision-based structure-from-motion models with accumulated point clouds from the LiDAR [11]. Although these approaches are generally more widely applicable as they enable sensor calibration from normal robot operation, they often still require an initial set of parameters.
@@ I Introduction
We propose to formulate extrinsic calibration as a graph optimization problem constrained by sensor motion and deep learning-based point correspondences.
@@ II Related Work
Correspondence-based methods replace the artificial targets with patterns that can be perceived in structured environments such as urban areas. For instance, SE-Calib [16] and Yin et al. [17] match object edges obtained from both images and LiDAR point clouds. Tu et al. [11] extract features of structure-from-motion (SfM) points of camera data and LiDAR points, followed by optimizing these correspondences jointly with camera intrinsics as well as camera and LiDAR poses. Koide et al. [18] build on the commonly used normalized information distance that poses a distance metric between the image and projected LiDAR points to measure the amount of mutual information. Finally, Caselitz et al. [19] propose a method for determining the pose of an RGB camera with respect to a 3D point cloud generated from LiDAR data by matching geometric clues. On the other hand, correspondence-free methods often rely on leveraging output data of auxiliary tasks such as monocular depth prediction [20, 21] or sensor motion estimation [10, 22, 17]. Both Zhang et al. [10] and Yin et al. [17] match trajectories from visual and LiDAR odometry and obtain extrinsic parameters via graph optimization. The latter then utilizes these parameters to initialize an edge-driven refinement stage. Finally, a more direct approach is enabled by exploiting deep learning-based correspondences between RGB images and LiDAR point clouds. Both RegNet [23] and CMRNext [24] involve training multiple CNNs on varying levels of decalibration and employ these networks during test time in a hierarchical manner. In detail, CMRNext frames the point-to-pixel matching problem as an optical flow estimation task. Building upon this approach, LCCNet [25] proposes the construction of a cost volume that stores matching costs. However, as it predicts the 6-DoF rigid-body transformation, it suffers from a high dependency on the training setup and is thus less generalizable.
@@ III-A Mathematical Preliminaries
We interpret camera-LiDAR calibration as an optimization problem that aims at determining the most likely transformation between the coordinate frames of a camera and a LiDAR given a set of sensor measurements. Mathematically, this can be represented by a conditional probability distribution:
@@
where denotes the robustifier and refers to a cost function applied to observation . For efficiency, we compute the squared Frobenius norm via vector multiplication of the induced error vectors with a diagonal weighting matrix . As we demonstrate in Sec. IV-C, the robustifier can significantly improve the optimization results by accounting for outliers in the observations.
To obtain the MAP estimate , we leverage graph optimization [15] with a single node representing the state and multiple constraints defined by the error terms . In particular, we define the state as the calibration parameters and decompose the sum into three error types.
with and observations for sensor motion and point correspondences, respectively.
@@ III-B Coarse Registration
where denotes the identity matrix. We further define the operator that reshapes a matrix to by stacking its columns. Note that we use the same information matrices and for all data pairs. In Fig. 4, we refer to the tuple of both error functions as odometry constraints. Finally, we perform MAP estimation to obtain the initial calibration parameters and scaling factors ***We omit these in the optimization equations to improve readability.:
@@ III-C Fine Registration
with denoting the information matrix for all . As indicated in Eq. 4, we do not add an error term for every image-point cloud pair. For instance, we do not apply for all but only to a subset reducing the number of partially redundant constraints in the optimization graph. Finally, we repeat the process of MAP estimation to obtain the overall calibration parameters :
@@ IV-A Robotic Platforms
We apply our method to a diverse set of sensor setups on four robotic platforms including the public KITTI dataset [31] as well as three in-house configurations.
@@
We provide results based on data from sequence 00 of the KITTI odometry benchmark [31]. Unlike many other works, we evaluate our approach and previous methods for both cameras. †: These methods did not release (English-speaking) code preventing reproducing results for both cameras.
KITTI Dataset: We extensively evaluate our method on the publicly available KITTI dataset [31]. In detail, we utilize the rectified images and undistorted point clouds of sequence 00 of the odometry dataset as data from the left camera of the other sequences was seen during the training of CMRNext [24]. The vehicle is equipped with two RGB cameras Point Grey Flea 2 and one LiDAR of type Velodyne HDL-64E. Sensor data is captured at .
@@ IV-B Main Results
For KITTI, we compare our MDPCalib approach to several classical [11, 17] and learning-based [20, 21, 24, 25, 32, 33] baselines in Sec. IV-A. If the authors released the corresponding code, we reproduce the calibration results for both the left and the right camera. Besides our MDPCalib, only the method by Yin et al. [17] is initialization-free. For the others, we follow the original sampling ranges. Note that the reproduced metrics of the learning-based approaches correspond to median calibration over multiple frames. For our method, we utilize 1,000 sensor motion constraints and 5% of the correspondences of 100 image-point cloud pairs. Although all baselines are outperformed by our approach, they generally yield accurate results for calibrating the LiDAR to the left camera. However, it is paramount to emphasize that all learning-based methods incorporate samples of the left camera in their training data. Therefore, we also calibrate the right camera measuring the capability to generalize. Except for CMRNext, all baselines suffer from a substantial performance drop. Nonetheless, MDPCalib yields the smallest error demonstrating that our joint optimization further increases robustness to unseen sensor configuration.
We confirm this observation by employing MDPCalib, CMRNext [24], and the method by Koide et al. [18] to three in-house robotic platforms. To account for the missing hardware time synchronization between camera and LiDAR as well as to alleviate inaccurate LiDAR undistortion, we only utilize static frames for computing point correspondences. As shown in Tab. II, MDPCalib yields significantly smaller calibration errors than previous methods. Given that we build on point correspondences from CMRNext, the results underline that our robust optimization scheme effectively ignores outliers in the point correspondences.
To foster future research, we report our calibration parameters for KITTI in the supplementary material. We show qualitative results represented by projecting the point clouds onto the images in Fig. 10 and provide further visualizations in the complementary video. Notably, the projections based on our calibration are still highly accurate at the maximum distance of .
@@ IV-C Ablation Studies
We extensively analyze the sensitivity of MDPCalib towards several design choices and adjustable parameters by conducting various ablation studies on the KITTI [31] dataset. In particular, we present results for calibrating the right RGB camera based on sequence 00. If not explicitly stated otherwise, we utilize 1,000 odometry poses and 5% of the point correspondences from 100 image-point cloud pairs. In Figs. 6, 8, 7 and 9, we visualize the mean and standard deviation of three runs. To improve readability, we add small offsets to the -values. For convenience, we provide the raw numerical metrics in the supplementary material.
Components Analysis: We analyze the impact of the calibration stages and the components of the optimization problem in Tab. III. Although the initialization-free coarse registration step based on sensor motion reduces the rotation error to sub-degree accuracy, the translation parameters suffer from a lack of observability. For reference, we provide the error of the median calibration parameters when running PnP on 100% of the predicted point correspondences. In Tab. III, we refer to this as CMRNext [24]. In MDPCalib, we instead process a subset of the correspondences as constraints in the graph optimization formulation, which further reduces the errors. In the bottom two rows, we demonstrate the efficacy of the key ingredients of MDPCalib. First, we show that utilizing a robust cost function significantly improves the results by addressing outliers in the observations. Second, while previous works [22, 17] only proposed to utilize sensor motion to initialize a subsequent correspondence-based calibration scheme, we incorporate both constraints in the refinement stage leveraging their complementary information. Our experiment clearly underlines the positive impact of joint optimization with respect to both sensor motion and point correspondences.
Runtime Analysis: We discuss the runtime of our method with respect to the number of both motion-based and point correspondences-based constraints. We conduct these experiments on a machine with an AMD Ryzen Threadripper PRO 3975WX CPU with and an NVIDIA A6000 GPU with . In Fig. 6 and Fig. 7, we visualize the runtime versus the number of poses and the relative number of correspondences, respectively. As can be seen in both experiments, the runtime scales approximately linearly, whereas the errors decrease. In both studies, we further observe an optimal configuration, after which the runtime continues to increase without major impacts on the accuracy.
@@
The last line highlighted in gray corresponds to our proposed MDPCalib. CMRNext denotes the errors of the median calibration parameters when running PnP on 100% of the predicted point correspondences after initialization with the coarse registration.
Number of Poses: In this study, we evaluate the effect of the number of poses from visual and LiDAR odometry on the calibration error, i.e., the length of the required data recording. In Fig. 6, we show the translation and rotation errors for an increasing number of poses. Although it generally holds that adding more motion constraints to the graph optimization results in more accurate calibration parameters, we observe that the error converges.
Number of Point Correspondences: Next, we repeat a similar study for the number of constraints induced by the point correspondences. For KITTI data, we measure an average of 20,000 correspondences per image-point cloud pair, i.e., our default setting of utilizing 5% generates approximately 1,000 constraints. In Fig. 7, we keep a fixed number of 100 pairs and vary the relative number of constraints. Fewer correspondences result in higher calibration errors. We confirm this observation in a second experiment visualized in Fig. 8, where we utilize 5% of the point correspondences but reduce the number of image-point cloud pairs. We note that in both experimental setups the errors converge towards a lower limit. Finally, we address the following question: Given a fixed number of correspondences, and hence runtime, is it preferable to increase the number of image-point cloud pairs or the relative amount of correspondences per pair? In Fig. 9, we plot the calibration errors for a stable number of point correspondences, i.e., when decreasing the number of pairs with respect to the default setting, we increase the relative number of correspondences by the respective amount. The experiment demonstrates that a smaller number of correspondences from a more diverse set of data pairs is preferable.
@@ V Conclusion
In this paper, we proposed MDPCalib, a novel method for the extrinsic calibration between camera and LiDAR sensors in an automated manner without the need for dedicated calibration targets. Our approach utilizes graph optimization to obtain the calibration parameters by aligning sensor motions from visual and LiDAR odometry and leveraging deep learning-based correspondences between 2D pixels and 3D points. In contrast to most previous learning-based methods, our approach generalizes to sensor configurations that differ from the training setup. Importantly, MDPCalib does not require accurate parameters for initialization. In practical experiments carried out on diverse robotic platforms, we demonstrated the efficacy of our method and provided detailed evaluations of several design choices. Future versions of our approach could incorporate constraints for multiple cameras and LiDARs and include intrinsic calibration in the optimization problem. Due to its unsupervised procedure, MDPCalib could further be extended to online calibration.

### v3

@@ Abstract
Sensor setups of robotic platforms commonly include both camera and LiDAR as they provide complementary information. However, fusing these two modalities typically requires a highly accurate calibration between them. In this paper, we propose MDPCalib which is a novel method for camera-LiDAR calibration that requires neither human supervision nor any specific target objects. Instead, we utilize sensor motion estimates from visual and LiDAR odometry as well as deep learning-based 2D-pixel-to-3D-point correspondences that are obtained without in-domain retraining. We represent camera-LiDAR calibration as an optimization problem and minimize the costs induced by constraints from sensor motion and point correspondences. In extensive experiments, we demonstrate that our approach yields highly accurate extrinsic calibration parameters and is robust to random initialization. Additionally, our approach generalizes to a wide range of sensor setups, which we demonstrate by employing it on various robotic platforms including a self-driving perception car, a quadruped robot, and a UAV. To make our calibration method publicly accessible, we release the code on our project website at http://calibration.cs.uni-freiburg.de.
Index Terms:
Calibration and Identification; Deep Learning Methods; Sensor FusionI Introduction
Sensor fusion for robotic systems has been extensively investigated [1, 2] as it promises to efficiently combine complementary information from different modalities, e.g., to increase robustness in case of sensor failures [3] and towards weather conditions [4]. However, the effectiveness of fusion approaches depends heavily on the extrinsic calibration between the sensors such as cameras and LiDAR.
Due to the importance of the task, camera-LiDAR calibration has been widely studied by the research community. Previously proposed methods can generally be classified into target-based and target-less approaches. Approaches of the first category often rely on artificial patterns such as checkerboards [5, 6] and require manual labor or another kind of human supervision to associate 2D points from the image space with 3D points in the LiDAR point cloud [7]. While a substantial effort has gone into the detection of the target and the automation of the matching process [8, 9], calibration often still needs special data collection. Some target-less calibration methods intend to overcome this problem, e.g., by inferring the extrinsic transform from the sensor motion [10] or by matching vision-based structure-from-motion models with accumulated point clouds from the LiDAR [11]. Although these approaches are generally more widely applicable as they enable sensor calibration from normal robot operation, they often still require an initial set of parameters.
@@
We propose to formulate extrinsic calibration as an optimization problem constrained by sensor motion and deep learning-based point correspondences.
@@ II Related Work
Correspondence-based methods replace the artificial targets with patterns that can be perceived in structured environments such as urban areas. For instance, Yuan et al. [16] and Yin et al. [17] match edges obtained from both images and LiDAR point clouds. Tu et al. [11] extract features of structure-from-motion (SfM) points of camera data and LiDAR points, followed by optimizing these correspondences jointly with camera intrinsics as well as camera and LiDAR poses. Koide et al. [18] build on the commonly used normalized information distance that poses a distance metric between the image and projected LiDAR points to measure the amount of mutual information. Finally, Caselitz et al. [19] propose a method for determining the pose of an RGB camera with respect to a 3D point cloud generated from LiDAR data by matching geometric clues. On the other hand, correspondence-free methods often rely on leveraging output data of auxiliary tasks such as monocular depth prediction [20, 21] or sensor motion estimation [10, 17, 22]. Both Zhang et al. [10] and Yin et al. [17] match trajectories from visual and LiDAR odometry and obtain extrinsic parameters via optimization. The latter then utilize these parameters to initialize an edge-driven refinement stage. Finally, a more direct approach is enabled by exploiting deep learning-based correspondences between RGB images and LiDAR point clouds. Both RegNet [23] and CMRNext [24] involve training multiple CNNs on varying levels of decalibration and employ these networks during test time in a hierarchical manner. In detail, CMRNext frames the point-to-pixel matching problem as an optical flow estimation task. Building upon this approach, LCCNet [25] proposes the construction of a cost volume that stores matching costs. However, as it predicts the 6-DoF rigid-body transformation, it suffers from a high dependency on the training setup and is thus less generalizable.
@@ III-A Mathematical Preliminaries
We interpret camera-LiDAR calibration as an optimization problem that determines the most likely transformation between the coordinate frames of a camera and a LiDAR given a set of sensor measurements. Mathematically, this can be represented by a conditional probability distribution:
@@
where denotes the robustifier and refers to a cost function applied to observation . For efficiency, we compute the squared Frobenius norm via vector multiplication of the induced error vectors with a diagonal weighting matrix . We utilize the Cauchy loss as robustifier due to its high tolerance to outliers in the observations [26], resulting in a significant increase in calibration accuracy.
To obtain the MAP estimate , we leverage an optimization formulation [15] with a single state and multiple constraints defined by the error terms . In particular, we define the state as the calibration parameters and decompose the sum into three error types:
with observations for sensor motion with rotational and translational costs and and observations for the point correspondences cost .
@@ III-B Coarse Registration
where denotes the identity matrix. We further define the operator that reshapes a matrix to by stacking its columns. Note that we use the same information matrices and for all data pairs. In Fig. 4, we refer to the tuple of both error functions as odometry constraints. Finally, we perform MAP estimation using the combined sensor motion cost to obtain the initial calibration parameters and scaling factors ***We omit these in the optimization equations to improve readability.:
@@ III-C Fine Registration
with denoting the information matrix for all . As indicated in Eq. 4, we do not add an error term for every image-point cloud pair. For instance, we do not apply for all but only to a subset reducing the number of partially redundant constraints in the optimization problem. Finally, we repeat the process of MAP estimation to obtain the overall calibration parameters :
@@ IV-A Robotic Platforms
We apply our method to a diverse set of sensor setups on four robotic platforms including the public KITTI dataset [32] as well as three in-house configurations. In the supplementary material, we provide further results on the Argoverse 2 [33] dataset.
@@
We provide results based on data from sequence 00 of the KITTI odometry benchmark [32]. Unlike many other works, we evaluate our approach and previous methods for both cameras. Bold and underlined values indicate the best and second-best scores, respectively. †: These methods did not release (English-speaking) code preventing reproducing results for both cameras.
KITTI Dataset: We extensively evaluate our method on the publicly available KITTI dataset [32]. In detail, we utilize the rectified images and undistorted point clouds of sequence 00 of the odometry benchmark as data from the left camera of the other sequences was seen during the training of CMRNext [24]. The vehicle is equipped with two RGB cameras Point Grey Flea 2 and one LiDAR of type Velodyne HDL-64E. Sensor data is captured at .
@@ IV-B Main Results
For KITTI, we compare our MDPCalib approach to several classical [11, 17] and learning-based [20, 21, 24, 25, 34, 35] baselines in Sec. IV-A. If the authors released the corresponding code, we reproduce the calibration results for both the left and the right camera. Besides our MDPCalib, only the method by Yin et al. [17] is initialization-free. For the others, we follow the original sampling ranges. Note that the reproduced metrics of the learning-based approaches correspond to median calibration over multiple frames. For our method, we utilize 1,000 sensor motion constraints and 5% randomly sampled correspondences from each of the 100 image-point cloud pairs. Although all baselines are outperformed by our approach, they generally yield accurate results for calibrating the LiDAR to the left camera. However, it is paramount to emphasize that all learning-based methods incorporate samples of the left camera in their training data. Therefore, we also calibrate the right camera measuring the capability to generalize. Except for CMRNext, all baselines suffer from a substantial performance drop. Nonetheless, MDPCalib yields the smallest error demonstrating that our joint optimization further increases robustness to unseen sensor configuration.
We confirm this observation by employing MDPCalib, CMRNext [24], and the method by Koide et al. [18] to three in-house robotic platforms. To account for the missing hardware time synchronization between camera and LiDAR, we only utilize static frames for computing point correspondences. As shown in Tab. II, MDPCalib yields significantly smaller calibration errors on the vehicle and the quadruped robot. For the UAV, we hypothesize that point cloud undistortion is more challenging due to less constrained motion resulting in less accurate odometry and hence larger errors.
We report our calibration parameters for KITTI in the supplementary material. We show qualitative results by projecting the point clouds onto the images in Fig. 10 and provide further visualizations in the complementary video. Notably, the projections based on our calibration are still highly accurate at the maximum distance of .
@@ IV-C Ablation Studies
We extensively analyze the sensitivity of MDPCalib towards several design choices and adjustable parameters by conducting various ablation studies on the KITTI dataset [32]. In particular, we present results for calibrating the right RGB camera based on sequence 00. If not explicitly stated otherwise, we utilize 1,000 odometry poses and 5% of the point correspondences from 100 image-point cloud pairs. In Figs. 6, 8, 7 and 9, we visualize the mean and standard deviation of three runs. To improve readability, we add small offsets to the -values. For convenience, we provide the raw numerical metrics in the supplementary material.
Components Analysis: We analyze the impact of the calibration stages and the components of the optimization problem in Tab. III. Although the initialization-free coarse registration step based on sensor motion reduces the rotation error to sub-degree accuracy, the translation parameters suffer from a lack of observability. For reference, we provide the error of the median calibration parameters when running PnP on 100% of the predicted point correspondences. In Tab. III, we refer to this as CMRNext [24]. In MDPCalib, we instead process a subset of the correspondences as constraints in the optimization formulation, which further reduces the errors. In the bottom two rows, we demonstrate the efficacy of the key ingredients of MDPCalib. First, we show that utilizing the Cauchy loss [26] to robustify the cost functions significantly improves the results by reducing the effect of outliers in the observations. Second, while previous works [22, 17] only proposed to utilize sensor motion to initialize a subsequent correspondence-based calibration scheme, we incorporate both constraints in the refinement stage leveraging their complementary information. Our experiment clearly underlines the positive impact of joint optimization with respect to both sensor motion and point correspondences.
Runtime Analysis: We discuss the runtime of our method with respect to the number of both motion-based and point correspondences-based constraints. We conduct these experiments on a machine with an AMD Ryzen Threadripper PRO 3975WX CPU with and an NVIDIA A6000 GPU with . In Fig. 6 and Fig. 7, we visualize the runtime versus the number of poses and the relative number of correspondences, respectively. As can be seen in both studies, the runtime scales approximately linearly, whereas the errors decrease. We further observe an optimal configuration, after which the runtime continues to increase without major impacts on the accuracy.
@@
The last line highlighted in gray corresponds to our proposed MDPCalib. CMRNext denotes the errors of the median calibration parameters when running PnP on 100% of the predicted point correspondences after initialization with the coarse registration. Bold and underlined values indicate the best and second-best scores, respectively.
Number of Poses: In this study, we evaluate the effect of the number of poses from visual and LiDAR odometry on the calibration error, i.e., the length of the required data recording. In Fig. 6, we show the translation and rotation errors for an increasing number of poses. Although it generally holds that adding more motion constraints to the optimization problem results in more accurate calibration parameters, we observe that the error converges.
Number of Point Correspondences: Next, we repeat a similar study for the number of constraints induced by the point correspondences. For KITTI data, we measure an average of 20,000 correspondences per image-point cloud pair, i.e., our default setting of utilizing 5% generates approximately 1,000 constraints. In Fig. 7, we keep a fixed number of 100 pairs and vary the relative number of constraints. Fewer correspondences result in higher calibration errors. We confirm this observation in a second experiment visualized in Fig. 8, where we utilize 5% of the point correspondences but reduce the number of image-point cloud pairs. Note that the errors converge towards a lower limit in both experimental setups. Finally, we address the following question: Given a fixed number of correspondences, and hence runtime, is it preferable to increase the number of image-point cloud pairs or the relative amount of correspondences per pair? In Fig. 9, we plot the calibration errors for a stable number of point correspondences, i.e., when decreasing the number of pairs with respect to the default setting, we increase the relative number of correspondences by the respective amount. The experiment shows that a smaller number of correspondences from a more diverse set of data pairs is preferable.
@@ V Conclusion
In this paper, we proposed MDPCalib, a novel method for the extrinsic calibration between camera and LiDAR sensors in an automated manner without the need for dedicated calibration targets. Our approach utilizes non-linear optimization to obtain the calibration parameters by aligning sensor motions from visual and LiDAR odometry and leveraging deep learning-based correspondences between 2D pixels and 3D points. In contrast to most previous learning-based methods, our approach generalizes to sensor configurations that differ from the training setup. Importantly, MDPCalib does not require accurate parameters for initialization. In practical experiments carried out on diverse robotic platforms, we demonstrated the efficacy of our method and provided detailed evaluations of several design choices. Future versions of our approach could incorporate constraints for multiple cameras and LiDARs and include intrinsic calibration in the optimization problem. Due to its unsupervised procedure, MDPCalib could further be extended to online calibration.